### v1

@@ 5 Conclusion
Automated video-EEG analysis holds significant potential for improving epilepsy diagnosis and treatment. Advances in deep learning and multimodal data integration have enhanced real-time seizure detection and prediction capabilities. However, challenges such as symptom variability, markup inconsistencies, and lack of publicly available data remain. Future research should focus on developing robust multimodal models, leveraging synthetic data, and creating standardized datasets to improve model performance and generalizability.
The proposed concept-based learning pipeline for treatment effect estimation offers a promising direction for integrating expert knowledge and multimodal data, ultimately advancing personalized epilepsy care.

### v2

@@ 5 Conclusion
Automated video-EEG analysis holds significant potential for improving epilepsy diagnosis and treatment. Advances in deep learning and multimodal data integration have enhanced real-time seizure and interictal epileptiform discharges detection and prediction capabilities. However, challenges such as symptom variability, markup inconsistencies, and lack of publicly available data remain. Future research should focus on developing robust multimodal models, leveraging synthetic data, and creating standardized datasets to improve model performance and generalizability.
The proposed concept-based learning pipeline for treatment effect estimation offers a promising direction for integrating expert knowledge and multimodal data, advancing personalized epilepsy care.
@@
Table 4: Papers on EEG-based epileptiform activity detection (2024 - Mar 2025)
Paper Title (†– open-source code) Citations Main Contribution
Detection of interictal epileptiform discharges using transformer based deep neural network for patients with self-limited epilepsy with centrotemporal spikes [tong_detection_2025] 2 Transformer-based IED detector. Evaulated on a private dataset and TUEV
Automatic Multi-label Classification of Interictal Epileptiform Discharges (IED) Detection Based on Scalp EEG and Transformer [rao_automatic_2024] 0 CNN + Transformer IED detector. Evaluated on a private dataset
Interictal Epileptiform Discharge Detection Using Time-Frequency Analysis and Transfer Learning [munia_interictal_2024] 0 ResNet-50 applied to TUEV
vEpiNet: A multimodal interictal epileptiform discharge detection method based on video and electroencephalogram data [lin_vepinet_2024] 3 Video (body and face detected with YOLOv5) + EEG (processed with EfficientNet2-S) for IED detection. Tested on a private dataset (vEpiSet is an EEG component of this dataset)
LightIED: Explainable AI with Light CNN for Interictal Epileptiform Discharge Detection [inoue_lightied_2024] 0 CNN for IED detection + explanation with Grad-CAM. Evaluated on a private dataset
Detection of Interictal epileptiform discharges with semi-supervised deep learning [de_sousa_detection_2024] 2 Semi-supervised VAE for IED detection. Evaluated on a private dataset
Robust compression and detection of epileptiform patterns in ECoG using a real-time spiking neural network hardware framework [costa_robust_2024] 16 IED and high frequency oscillation (HFO) detector on a low power consuming chip. Dataset made publicly available
Advancing epilepsy diagnosis: A meta-analysis of artificial intelligence approaches for interictal epileptiform discharge detection [borges_camargo_diniz_advancing_2024] 1 An analysis of 23 studies (IED detection models)
Method for cycle detection in sparse, irregularly sampled, long-term neuro-behavioral timeseries: Basis pursuit denoising with polynomial detrending of long-term, inter-ictal epileptiform activity [balzekas_method_2024] 2 A method for IED rate analysis (IED detection not covered)
A review of signal processing and machine learning techniques for interictal epileptiform discharge detection [abdi-sargezeh_review_2024] 14 A review of IED detection approaches